OpenStudy (anonymous):

(2^(-2)/x^4)^(-2) * ((-2)x^3/x-1)^3

myininaya (myininaya):

\[(2^{-2}x^{-4})^{-2}\cdot(\frac{-2x^3}{x-1})^3\]

myininaya (myininaya):

\[(2^{-2})^{-2}(x^{-4})^{-2}\frac{(-2)^{3}(x^3)^3}{(x-1)^3}\]

myininaya (myininaya):

\[2^4x^8\frac{-8x^9}{(x-1)^3}\]

myininaya (myininaya):

\[16x^8\frac{-8x^9}{(x-1)^3}\] \[\frac{16(-8)x^8x^9}{(x-1)^3}\]

myininaya (myininaya):

\[\frac{-128x^{17}}{(x-1)^3}\]

myininaya (myininaya):

hopefully i read the problem as you meant it

OpenStudy (anonymous):

you did myininaya, thnx

myininaya (myininaya):

:)